OpenStudy (anonymous):

Holly is taking a multiple choice reading test. She has decided to guess on questions 3 and 7. On each question, there are four worded choices plus an "all of these" choice and a "none of these" choice. What is the probability that she answers question 3 incorrectly and question 7 correctly?

OpenStudy (anonymous):

hmm...it seems there is 4 answers plus 2 other which means that both 3 and 7, the chance of getting them correct would be 1/6. however, you want to find the chances of getting on wrong and the next right so.. question's 3 chances of getting it wrong= 5/6 question's 7 chances of getting it right = 1/6 you must multiple both together to get the number which would be 5/6 x 1/6 = 5/36= 13.88% Hope this helps